OpenStudy (anonymous):

A copy machine makes 44 copies per minute. How many copies does it make in 4 minutes and 15 seconds?

OpenStudy (anonymous):

Every minute it makes 44 copies so in 4 minutes it will make 4*44 = 176. 15 seconds is 1/4 of a minute so in that time it can make 1/4 * 44 = 11. In total it will make 176 + 11 = 187